Calculate the pH of a buffer solution made by mixing 45.0 ml of 0.25 M CH3COOH and 25.0 ml of 0.15 M solution of CH3COONa.
1
Expert's answer
2020-05-13T14:19:48-0400

V(buffer solution)=0,045+0,025=0,07

C1*V1=C2*V2 => C(CH3COOH in buffer solution)=(C1*V1)/V2=(0,045*0,25)/(0,045+0,025)=0,161

C(CH3COONa in buffer solution)=(C1*V1)/V2=(0,025*0,15)/(0,045+0,025)=0,0536

pH=pK(CH3COOH)+lg(C(CH3COONa)/C(CH3COOH))= 4,27